# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

### v1

@@ 4.4. The optimal control problem
Given Theorems 4.10 and 4.11, the proof of the following dynamic programming principle is standard (see, e.g., Proposition III.2.5 in [BardiCapuzzoDolcetta]).
@@
Let $(s,x)\in(0,T]\times\Omega$ and $s_{1}\in[s,T]$ Then
@@
4.5. Proofs of Theorems 4.10 and  4.11
We will make use of an equivalent path-dependent discrete-time infinite-horizon decision model. (For corresponding work in the Markovian case, see, e.g., section 3 of [BR10PDMP], section 8.2 in [BR11book], and [RTT17PDMP].) In this section, fix $(s,x)\in(0,T]\times\Omega$.
First, we determine appropriate data for our discrete-time model. Let $\mathbf{a}\in\mathbb{A}^{\prime}$. We have
$J(s,x,\mathbf{a})$ $=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Bigl{[}\int_{T_{n}}^{T_{n+1}}{\mathbf{1}}_{[0,T]}(t)\,\ell(t,X^{s,x},\mathcal{A}_{n}(t))\,{\mathrm{d}}t+{\mathbf{1}}_{[T_{n},T_{n+1})}(T)\,h(X^{s,x})\Bigr{]}.$
${\mathbf{1}}_{[0,T]}(t_{n+1})\,h(X^{s,x}(\check{\omega}))={\mathbf{1}}_{[0,T]}(t_{n+1})\,h(X^{s,x}_{n}(\check{\omega}_{n+1}))$
$\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Bigl{[}\int_{T_{n}}^{T_{n+1}}{\mathbf{1}}_{[0,T]}(t)\,\ell(t,X^{s,x},\mathcal{A}_{n}(t))\,{\mathrm{d}}t+{\mathbf{1}}_{[T_{n},T_{n+1})}(T)\,h(X^{s,x})\Bigr{]}$
$=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{n+2}\Bigl{[}\int_{T_{n}\wedge T}^{T}{\mathbf{1}}_{(t,\infty]}(T_{n+1})\ell(t,X_{n}^{s,x},\mathcal{A}_{n}(t))\,{\mathrm{d}}t+{\mathbf{1}}_{[0,T]}(T_{n}){\mathbf{1}}_{(T,\infty]}(T_{n+1})\,h(X_{n}^{s,x})\Bigr{]}$
$=\int_{\check{\Omega}_{n+1}}\Big{(}\int_{t_{n}\wedge T}^{T}\chi^{t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(t)\,\ell(t,X^{s,x}_{n}(\check{\omega}_{n+1}),a_{n}(t))\,{\mathrm{d}}t$
Consequently,
(4.8) $J(s,x,\mathbf{a})=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}_{n+1}^{s,x,\mathbf{a}}\left[\vec{\ell}_{n}\right],$
where the functions $\vec{\ell}_{n}:\check{\Omega}_{n+1}\to{\mathbb{R}}_{+}$, $n\in{\mathbb{N}}_{0}$, are defined by
For each $n\in{\mathbb{N}}_{0}$, define a stochastic kernel $\vec{Q}_{n}$ on (+×$({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}$ given ((+×$(({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\})^{n+1}\times\mathcal{A}^{n+1}$ by
(+×$\int_{({\mathbb{R}}_{+}\times E)\cup\{\infty,\Delta\}}{\mathbf{1}}_{[0,\infty)}(t)\,\psi(t,e)\,\vec{Q}_{n}({\mathrm{d}}t\,{\mathrm{d}}e|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n})$
:={+×$:=\begin{cases}&-\int_{{\mathbb{R}}_{+}\times E}\psi(t,e)\,Q(t,{\mathrm{d}}e|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\,{\mathrm{d}}F(t|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\\ &\qquad\qquad\text{if $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}\in\check{\Omega}_{n+1}$,}\\ &\psi(\infty,\Delta)\qquad\text{otherwise,}\end{cases}$
for every bounded Borel-measurable function $\psi:{\mathbb{R}}_{+}\times E\cup\{(\infty,\Delta)\}\to{\mathbb{R}}$.
Remark 4.13.
Borel-measurability of $\vec{Q}_{n}$ follows from Remark 4.3 and Remark 4.6. Keep in mind that $(s,x)$ is fixed here (see the beginning of section 4.5).
Next, having determined the mentioned data, we put them into the context of section A, which is devoted to discrete-time stochastic control problems. This will allow us to use the (discrete-time) Bellman equation (A.7), which in turn will yield (4.6).
Model 4.14.
Consider a discrete-time infinite-horizon stochastic optimal control model from section A.1 with the following data:
  • State space: $S=({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}$.
  • Control space: $U=\mathcal{A}\cup\{\Delta^{\prime}\}$.
  • Disturbance kernel $p_{k}=\vec{Q}_{k}$.
  • Cost function $g_{k}=\vec{\ell}_{k}.{\mathbf{1}}_{\check{\Omega}_{k+1}}+\infty.{\mathbf{1}}_{(S^{k+1}\times U^{k+1})\setminus\check{\Omega}_{k+1}}$ in the sense that
    $g_{k}((t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n}):=\begin{cases}\vec{\ell}_{k}((t_{j},e_{j},\alpha_{j})_{j=0}^{n})\text{ if $(t_{j},e_{j},\alpha_{j})_{j=0}^{n}\in\check{\Omega}_{n+1}$,}\\ \infty\text{ otherwise.}\end{cases}$
With the just specified data, the corresponding marginals $r_{N}(\pi,p)$ on $(S\times U)^{N}$, $N\in{\mathbb{N}}$ defined by (A.1) satisfy
$\int g_{0}\,{\mathrm{d}}r_{1}\left(\mathbf{a},\delta_{(s,x(s)}\right)$ $=\int_{S\setminus(\{0\}\times E)}\int_{U}\vec{\ell}_{0}(t_{0},e_{0},\alpha_{0})\,\mathbf{a}_{0}({\mathrm{d}}\alpha_{0})\,\delta_{(s,x(s))}({\mathrm{d}}t_{0}\,{\mathrm{d}}e_{0})$
+.$\qquad+\infty.\delta_{(s,x(s))}(\{0\}\times E)\qquad\text{(note that $s>0$)}$
$=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{1}\left[\vec{\ell}_{0}\right],$
$\int g_{1}\,{\mathrm{d}}r_{2}\left(\mathbf{a},\delta_{(s,x(s))}\right)$ $=\int_{S}\int_{U}\int_{S}\int_{U}g_{1}((t_{0},e_{0}),(t_{1},e_{1});\alpha_{0},\alpha_{1})$
$=\check{{{\mathbb{E}}}}^{s,\mathbf{a}}_{2}\left[\vec{\ell}_{1}\right]\qquad\text{because $\left[r_{2}\left(\mathbf{a},\delta_{(s,x(s))}\right)\right](\check{\Omega}_{2})=1$,}$
[…]
Thus, together with the definition of the optimal cost $J^{\ast}$ in (A.3), we have
(4.10) $V^{\prime}(s,x)$ $=J^{\ast}(s,x(s)).$
We need the next statement. Its proof is postponed in section B.2.
Lemma 4.15.
Fix $(s,x)\in(0,T)\times\Omega$ and $(t,e,\alpha)\in(s,T]\times E\times\mathcal{A}$. Then
$V^{\prime}(t,\phi^{s,x,\alpha}\otimes_{t}e)=J^{\ast}(1;(s,x(s)),(t,e);\alpha),$
where $J^{\ast}(1;\cdot)$ is defined by (A.6) with the data of Model 4.14.
By Theorem A.3 (below the first $\Delta^{\prime}$ can be omitted because we have $\alpha_{0}=\Delta^{\prime}$ if and only if $t_{0}=\infty$ but $t_{0}=s\in(0,T]$) and by noting that the running cost vanishes if $t>T$ due to (4.9), we have
$J^{\ast}(s,x(s))=\inf_{\alpha\in\mathcal{A}\cup\{\Delta^{\prime}\}}\Bigl{\{}g_{0}((s,x(s));\alpha)$
+(+×$\qquad\qquad+\int_{({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}}J^{\ast}(1;(s,x(s)),(t,e);\alpha)\,\vec{Q}_{0}({\mathrm{d}}t\,{\mathrm{d}}e|(s,x(s));\alpha)\Bigr{\}}$
$=\inf_{\alpha\in\mathcal{A}}\Bigl{\{}g_{0}((s,x(s));\alpha)$
Together with (4.10) and Lemma 4.15, we get (4.6). This concludes the proof of Theorem 4.10.
By Theorem A.4,
$J^{\ast}(s,x(s))=\inf_{\mathbf{a}\in\mathbb{A}}J_{\mathbf{a}}(s,x(s)),$
where $J_{\mathbf{a}}$ is defined by (A.2). Thus
$V^{\prime}(s,x)=V(s,x).$
This concludes the proof of Theorem 4.11. ∎
5. Non-local path-dependent Hamilton-Jacobi-Bellman equations
@@
In this section, the following additional assumption is in force. This will allow us to use the results of section 7 in [BK1].
@@
We study a terminal-value problem with a non-local path-dependent partial differential equation of the form
(5.1) $\begin{split}-\partial_{t}u(t,x)-F(t,x,u(t,x),\partial_{x}u(t,x),u(t,x\otimes_{t}\cdot))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega,\end{split}$
@@
where $F:[0,T]\times\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}\times B(E)\to{\mathbb{R}}$ is defined by
$F(t,x,y,z,\psi)$ $:=\operatornamewithlimits{essinf}_{a\in A}\Bigl{\{}\ell(t,x,a)+(f(t,x,a),z)$
@@
$\qquad\qquad+\lambda(t,x,a)\,\int_{{\mathbb{R}}^{d}}\left[\psi(e)-y\right]\,Q(t,x,a,{\mathrm{d}}e)\Bigr{\}}.$
We aim at proving the following result (see Definition 5.10 below for our notion of minimax solution):
@@
The operator $G$ defined by (5.2) has a unique fixed point $\bar{\psi}$, which is the unique minimax solution to (5.1) and coincides with the value function $V$ defined by (4.5) on $(0,T]\times\Omega$.
We continue as follows: First, we introduce the mentioned operator $G$, study its properties, and the properties of its fixed point. Then, we carry out the proof of Theorem 5.2. Finally, we establish a comparison principle between minimax sub- and supersolutions to (5.1).
5.1. The operator $G$
Let $\mathrm{BLSA}([0,T]\times\Omega)$ be the space of all bounded lower semi-analytic functions from $[0,T]\times\Omega$ to ${\mathbb{R}}$.
Given $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$ and $(s,x)\in[0,T]\times\Omega$, define
(5.2) $\begin{split}&[G\psi](s,x):=\inf_{a\in\mathcal{A}}\Bigl{\{}\chi^{s,x,a}(T)\,h(\phi^{s,x,a})\\ &+\int_{s}^{T}\chi^{s,x,a}(t)\,\Bigl{[}\ell^{s,x,a}(t)+\int_{{\mathbb{R}}^{d}}\psi(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t\Bigr{\}}.\end{split}$
Remark 5.3.
By Theorem 96, 146-IV, in [DMprobPot], Lemma 4.4, Remark 4.5, Proposition 7.48 in [BS], and Proposition 7.47 in [BS],
(5.3) $G:\mathrm{BLSA}([0,T]\times\Omega)\to\mathrm{BLSA}([0,T]\times\Omega).$
Let us introduce several operators related to $G$.
Given $s_{1}$, $s_{2}\in{\mathbb{R}}$ with $0\leq s_{1}<s_{2}\leq T$, $\eta\in\mathrm{BLSA}([s_{2},T]\times\Omega)$, and $a\in\mathcal{A}$, define131313 This is possible for the same reason that (5.3) in Remark 5.3 holds.
(5.4) $\begin{split}G_{s_{1},s_{2};\eta,a}&:\mathrm{BLSA}([s_{1},s_{2}]\times\Omega)\to\mathrm{BLSA}([s_{1},s_{2}]\times\Omega),\,\psi\mapsto(G_{s_{1},s_{2};\eta,a})\psi,\\ G_{s_{1},s_{2};\eta}&:\mathrm{BLSA}([s_{1},s_{2}]\times\Omega)\to\mathrm{BLSA}([s_{1},s_{2}]\times\Omega),\,\psi\mapsto(G_{s_{1},s_{2};\eta})\psi,\end{split}$
by
$[(G_{s_{1},s_{2};\eta,a})\psi](s,x):=\chi^{s,x,a}(s_{2})\,\eta(s_{2},\phi^{s,x,a})$
$\qquad+\int_{s}^{s_{2}}\Bigl{[}\chi^{s,x,a}(t)\,\ell^{s,x,a}(t)+\int_{{\mathbb{R}}^{d}}\psi(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\chi^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t$
and
$[(G_{s_{1},s_{2};\eta})\psi](s,x)$ $:=\inf_{\tilde{a}\in\mathcal{A}}[(G_{s_{1},s_{2};\eta,\tilde{a}})\psi](s,x)$
for every $(s,x)\in[s_{1},s_{2}]\times\Omega$.
Theorem 5.4.
The following holds:
(i) The operator $G$ defined by (5.2) has a unique fixed point $\bar{\psi}$ in $\mathrm{BLSA}([0,T]\times\Omega)$, which coincides with $V$ on $(0,T]\times\Omega$.

### v2

@@ 4.4. A related discrete-time optimal control problem and the dynamic programming principle
Proof.
By Theorem 4.16 (below the first $\Delta^{\prime}$ can be omitted because we have $\alpha_{0}=\Delta^{\prime}$ if and only if $t_{0}=\infty$ but $t_{0}=s\in(0,T]$) and by noting that the running cost vanishes if $t>T$ due to (4.7), we have
$J^{\ast}(0;s,x(s))=\inf_{\alpha\in\mathcal{A}\cup\{\Delta^{\prime}\}}\Bigl\{g_{0}((s,x(s));\alpha)$
$\qquad\qquad+\int_{({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}}J^{\ast}(1;(s,x(s)),(t,e);\alpha)\,\vec{Q}_{0}({\mathrm{d}}t\,{\mathrm{d}}e|(s,x(s));\alpha)\Bigr\}$
$=\inf_{\alpha\in\mathcal{A}}\Bigl\{g_{0}((s,x(s));\alpha)$
Together with Theorem 4.15, we get (4.18). ∎
For the next theorem, recall the set $\mathbb{A}$ of non-randomized policies from Definition 4.9.
Theorem 4.18.
Let $(t_{0},e_{0})\in S$. Then $J^{\ast}(0;t_{0},e_{0})=\inf_{\mathbf{a}\in\mathbb{A}}J_{\mathbf{a}}(0;t_{0},e_{0})$.
Proof.
This result is a specific instance of Theorem A.4, whose proof is provided in appendix A. ∎
@@
Proof.
Let $(s,x)\in(0,T]\times\Omega$. By Theorem 4.18, $J^{\ast}(0;s,x(s))=\inf_{\mathbf{a}\in\mathbb{A}}J_{\mathbf{a}}(0;s,x(s))$. Together with Theorem 4.15 and (4.14), we can deduce that $V^{\prime}(s,x)=V(s,x)$. ∎
Given Theorems 4.17 and 4.19, the proof of the following dynamic programming principle is standard (see, e.g., Proposition III.2.5 in [BardiCapuzzoDolcetta]).
@@
Let $(s,x)\in(0,T]\times\Omega$ and $s_{1}\in[s,T]$ . Then
@@
5. Non-local path-dependent Hamilton–Jacobi–Bellman equations
@@
In this section, the following additional assumption is in force. This will allow us to use the results of section 7 in our companion paper [BK1].
@@
We study the terminal-value problem
(5.1) $\begin{split}-\partial_{t}u(t,x)-{\color[rgb]{0,0,0}H^{u}(t,x,u(t,x),\partial_{x}u(t,x))}&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega,\end{split}$
@@
where $(t,x,y,z,\psi)\mapsto H^{\psi}(t,x,y,z)$, $0,T]\times\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}\times\mathrm{BLSA}([0,T]\times\Omega)\to{\mathbb{R}}$, is defined by
@@
(5.2) $\begin{split}{\color[rgb]{0,0,0}H^{\psi}(t,x,y,z)}&:={\color[rgb]{0,0,0}\inf_{a\in A}}\Bigl\{\ell(t,x,a)+(f(t,x,a),z)\\ &\qquad\qquad+\lambda(t,x,a)\,\int_{{\mathbb{R}}^{d}}\left[\psi(t,x\otimes_{t}e)-y\right]\,Q(t,x,a,{\mathrm{d}}e)\Bigr\}.\end{split}$
Remark 5.2.
Recall from section 3.1 that $0,T]\times\Omega$ and its subsets are considered to be equipped with the pseudo-metric defined in (3.1). This implies that a function $u:[0,T]\times\Omega\to{\mathbb{R}}$ that is semicontinuous (with respect to this pseudo-metric) is non-anticipating, i.e., $u(t,x)=u(t,x_{\cdot\wedge t})$ for each $(t,x)\in[0,T]\times\Omega$.
Let us now outline our approach for establishing existence and uniqueness of (5.1). First note that (5.1) formally corresponds to the HJB equation associated to the deterministic optimal control problem that is specified by the value function
(5.3) $v(s,x)=\inf_{\alpha\in\mathcal{A}}\Bigl\{\chi^{s,x,\alpha}(T)\,h(\phi^{s,x,\alpha})+\int_{s}^{T}\chi^{s,x,\alpha}(t)\,\ell^{v}(t,\phi^{s,x,\alpha},\alpha(t))\,{\mathrm{d}}t\Bigr\}$
where the running cost
$\ell^{v}(t,x,a):=\ell(t,x,a)+\int_{{\mathbb{R}}^{d}}v(t,x\otimes_{t}e)\lambda(t,x,a)\,Q(t,x,a,{\mathrm{d}}e)$
depends on the value function $v$ itself (cf. Theorem 4.17). To tackle this issue related to the running cost, we consider related deterministic optimal control problems specified by value functions $v^{\psi}:[0,T]\times\Omega\to{\mathbb{R}}$ parametrized by functions $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$. These value functions are defined as follows:
(5.4) $v^{\psi}(s,x):=\inf_{\alpha\in\mathcal{A}}\Bigl\{\chi^{s,x,\alpha}(T)\,h(\phi^{s,x,\alpha})+\int_{s}^{T}\chi^{s,x,\alpha}(t)\,\ell^{\psi}(t,\phi^{s,x,\alpha},\alpha(t))\,{\mathrm{d}}t\Bigr\}.$
Here, the running costs $\ell^{\psi}:[0,T]\times\Omega\to A$ are given by
$\ell^{\psi}(t,x,a):=\ell(t,x,a)+\int_{{\mathbb{R}}^{d}}\psi(t,x\otimes_{t}e)\lambda(t,x,a)\,Q(t,x,a,{\mathrm{d}}e).$
Each $v^{\psi}$ is expected to be a unique solution of the terminal value problem
($\text{{TVP}}_{\psi}$) $\begin{split}-\partial_{t}u(t,x)-{\color[rgb]{0,0,0}H^{\psi}(t,x,u(t,x),\partial_{x}u(t,x))}&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega.\end{split},$
This topic is addressed in Theorem 5.8 at the end of subsection 5.2. Next, we establish in subsection 5.3 that the operator [^5] See (5.4) for the definition of $v^{\psi}$.
(5.6) $G:\psi\mapsto v^{\psi},\,\mathrm{BLSA}([0,T]\times\Omega)\to{\mathbb{R}}^{[0,T]\times\Omega},$
has a unique fixed point $\bar{\psi}$ in $\mathrm{BLSA}([0,T]\times\Omega)$, which satisfies (5.3) in place of $v$. This result together with existence and uniqueness for ($\text{{TVP}}_{\psi}$) will fairly quickly lead to existence and uniqueness for minimax solutions of (5.1) in subsection 5.4. Also note that $\bar{\psi}$ coincides on $(0,T]\times\Omega$ with the value function $V$ defined in (4.6) of our stochastic optimal control problem treated in section 4, i.e.,
$-\partial_{t}u(t,x)-{\color[rgb]{0,0,0}H^{u}(t,x,u(t,x),\partial_{x}u(t,x))}=0$
is the HJB equation associated to the stochastic optimal control problem for path-dependent PDPs specified by (4.6). Finally, we will obtain a comparison principle between minimax sub- and and supersolutions of (5.1) in subsection 5.5.
The next subsections 5.1 and 5.2 are devoted to the introduction of basic material such as path derivatives for functions on $0,T]\times\Omega$ and notions of solutions for (5.1).
5.1. Path derivatives and classical solutions
The concept of derivatives on $0,T]\times\Omega$ that we are using here is a slight modification of the notion of coinvariant derivatives due to A. V. Kim [Kim85].
Definition 5.3.
The space $C_{b}^{1,1}([0,T]\times\Omega)$ is the set of all $u\in C_{b}([0,T]\times\Omega)$, for which there exist functions $\partial_{t}u\in C_{b}([0,T]\times\Omega)$ and $\partial_{x}u\in C_{b}([0,T],{\mathbb{R}}^{d})$ also called path derivatives of $u$ such that, for every $(t_{0},x_{0})\in[0,T)\times\Omega$, every $x\in\Omega$ that is Lipschitz continuous on $t_{0},T]$ and coincides with $x_{0}$ on $0,t_{0}]$, and every $t\in(t_{0},T]$,
$u(t,x)-u(t_{0},x_{0})=\int_{t_{0}}^{t}(\partial_{t}u(s,x)+\partial_{x}u(s,x)\cdot x^{\prime}(s))\,{\mathrm{d}}s.$
Note that if $u\in C^{1,1}_{b}([0,T]\times\Omega)$, then its path derivatives $\partial_{t}u$ and $\partial_{x}u$ are uniquely determined. Moreover, there is a close relationship of the here defined path derivatives with Dupire’s explicitly defined horizontal and vertical path derivatives introduced in [dupirefunctional]. We refer to section 3 of our companion paper [BK1] for details.
Definition 5.4.
A function $u:[0,T]\times\Omega\to{\mathbb{R}}$ is a classical solution of (5.1) if $u\in C_{b}^{1,1}([0,T]\times\Omega)$, $u(T,\cdot)=h$, and, for every $(t,x)\in[0,T)\times\Omega$,
$-\partial_{t}u(t,x)-{\color[rgb]{0,0,0}H^{u}(t,x,u(t,x),\partial_{x}u(t,x))}=0.$
5.2. Minimax solutions
Here, we define our notion of nonsmooth solutions of the terminal value problems (5.1) and ($\text{{TVP}}_{\psi}$), $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$. To this end we need the following path spaces. Given $(s_{0},x_{0})\in[0,T)\times\Omega$ and $z\in{\mathbb{R}}^{d}$, define
$\mathcal{X}(s_{0},x_{0})$ $:=\Big\{x\in\Omega:\,x=x_{0}\text{ on $[0,s_{0}]$, $x|_{[s_{0},T]}$ is absolutely continuous with}$
$\qquad\qquad\left|{x^{\prime}(t)}\right|\leq C_{f}(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ a.e.~on $(s_{0},T)$}\Big\}.$
Given $(s_{0},x_{0})\in[0,T)\times\Omega$, $z\in{\mathbb{R}}^{d}$, and $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$, define
(5.7) $\begin{split}&\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z):=\Big\{(x,y)\in\mathcal{X}(s_{0},x_{0})\times C([s_{0},T]):\\ &\qquad y(t)=y_{0}+\int_{s_{0}}^{t}(x^{\prime}(s),z)-{\color[rgb]{0,0,0}H^{\psi}(s,x,y(s),z)}\,{\mathrm{d}}s\text{ on $[s_{0},T]$}\Big\}.\end{split}$
Definition 5.5.
Let $u:[0,T]\times\Omega\to{\mathbb{R}}$.
(i) $u$ is a minimax supersolution of (5.1) if $u\in{\color[rgb]{0,0,0}\mathrm{BLSC}}([0,T]\times\Omega)$, if $u=u(s,x)$ is Lipschitz in $x$, if $u(T,\cdot)\geq h$, and if, for all $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$ and all $y_{0}\geq u(s_{0},x_{0})$, there is an $(x,y)\in\mathcal{Y}_{u}(s_{0},x_{0},y_{0},z)$ such that $y(t)\geq u(t,x)$ for each $t\in[s_{0},T]$.
(i) $u$ is a minimax subsolution of (5.1) if $u\in{\color[rgb]{0,0,0}\mathrm{BUSC}}([0,T]\times\Omega)$, if $u=u(s,x)$ is Lipschitz in $x$, if $u(T,\cdot)\leq h$, and if for all $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$ and all $y_{0}\leq u(s_{0},x_{0})$, there is an $(x,y)\in\mathcal{Y}_{u}(s_{0},x_{0},y_{0},z)$ such that $y(t)\leq u(t,x)$ for each $t\in[s_{0},T]$.
(iii) $u$ is a minimax solution of (5.1) if $u$ is both a minimax supersolution and a minimax subsolution of (5.1).
@@ 5.2. Minimax solutions
Let $u\in C_{b}^{1,1}([0,T]\times\Omega)$. Suppose in addition that $u=u(t,x)$ is Lipschitz in $x$ and that all mappings
(5.8) $(t,x)\mapsto{\color[rgb]{0,0,0}H^{u}(t,x,u(t,x),z)},\quad[0,T]\times\Omega\to{\mathbb{R}},\quad z\in{\mathbb{R}}^{d},$
are continuous. Then $u$ is a classical solution of (5.1) if and only if $u$ is a minimax solution of (5.1).